# DoubleAgents: Exploring Mechanisms of Building Trust with Proactive AI

### v1

@@ Abstract.
Agentic workflows promise efficiency, but adoption hinges on whether people actually trust systems that act on their behalf. We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks. A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their reliance before live use. We evaluate DoubleAgents in a two-day lab study (n=10), two deployments (n=2), and a technical evaluation. Results show that participants initially hesitated to delegate but grew more reliant as they experienced transparency, control, and adaptive learning during simulated cases. Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that the effort required scaled appropriately with task complexity and contextual data. We contribute trust-by-design patterns and mechanisms for proactive AI—consistency, controllability, and explainability—along with simulation as a safe path to build and calibrate trust over time.
@@
copyright: acmcopyrightjournalyear: 2025conference: ; ; booktitle: Proceedings of the 2025price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
Coordinating human effort is essential for tasks at every scale. Event planning involves finding suitable times, locations, vendors, and transportation. Warehouses must assign workers with the right skills and experience to shifts they are available for. Even organizing a speaker series requires emailing potential speakers, collecting their availability, and assigning them to appropriate slots. These problems involve both discovering people’s availability and solving the constraints of assigning individuals to limited roles. Human coordinators face time-consuming and intricate communication challenges: people may hesitate to disclose their full availability, delay responses, or provide only tentative answers such as “I might be able to come Saturday, but I’m not sure yet.” Coordinators must also balance group dynamics, make exceptions for important individuals, and continuously replan when last-minute changes arise. Coordination is thus both tedious and delicate, requiring continuous monitoring and re-planning while respecting social norms and relationships.
Generative AI now possesses several key capabilities that could help with coordination, including deep reasoning to evaluate constraints and produce plans (Guo et al., 2025; Yao et al., 2023b, a; Stechly et al., 2025), and API tool-calling to take concrete actions (Trivedi et al., 2024; Yao et al., 2024; Masterman et al., 2024). However, people do not yet trust AI to act proactively on their behalf for complex and delicate tasks like human coordination problems. AI systems are prone to hallucinations, often lack sufficient context to understand the problem (Zhang et al., 2025; Athaluri et al., 2023), and may mishandle delicate social interactions (Zhou et al., 2023; Wang et al., 2024). Most importantly, they lack an awareness of when to stop and escalate to human oversight. As a result, current AI interfaces largely wait for users to issue prompts and micromanage responses. This limits productivity and prevents AI from relieving people of burdensome procedural work. To unlock proactive AI, we need frameworks that let humans observe, guide, and calibrate trust in AI over time.
Trust is not established by statements alone—it develops through repeated reliable behavior over time (Razin and Feigh, 2024; Fouragnan et al., 2013). Game theory suggests that trust between humans is built by repeated observation of behavior in varied circumstances (Kollock, 1998). Additionally, in the AI trust and safety literature, Constitutional AI (Bai et al., 2022) is a paradigm that uses human-articulated policies (called a constitution) that can align an AI’s behavior with human expectations and values. Policies in making expectations explicit - so that humans and AI have a shared understanding of the expected behavior. Repeated interactions and clear policies are two pillars of establishing trust that we build upon to help build trust in proactive AI.
@@
We propose to build trust in proactive AI through simulation and policy refinement. Human behavior can be unpredictable, but AI-driven simulations can generate a wide range of realistic scenarios (Park et al., 2023), including late, incomplete, or uncertain replies. These allow coordinators to test and refine policies in safe environments. Policies specify how the AI should act (e.g., when to follow up, when to stop, how to phrase outreach) and when to escalate to human oversight. Hence, we propose to use simulation to expose AI behavior across repeated interactions, letting users correct and evolve policies iteratively. Visualizations of system state—goal progress, a timeline view, and person-level interaction histories—provide transparency and help humans intervene efficiently when needed.
We present DoubleAgents, shown in Figure 1, an interactive agentic system designed to automate aspects of human coordination while supporting user oversight and trust-building. DoubleAgents integrates generative AI agents in two complementary roles: (1) a ReAct Loop (Yao et al., 2023b) with AI planning and execution agents that can call tools to get appropriate context, query the current state of the system, and execute actions such as emailing potential speakers. This loop is extended to provide interaction for human interactivity and oversight at every step. (2) A LLM-based simulation of respondents that generate the varied and realistic human responses for various respondent personas (e.g. a professor who is very slow to respond, or a speaker who won’t commit). The LLM simulations allow users to experience repeated execution of the system to understand how it works, evolve the policies over time to suit their preferences, and potentially build trust that the system will act according to their wishes. By combining policy generation and refinement, action execution, and multi-agent LLM simulation within a user-centered feedback process, DoubleAgents helps users delegate repetitive aspects of coordination while retaining control of high-stakes interactions.
In a lab study (n=10) of people using the system with simulated responses over two study sessions, participants began hesitant to delegate but, through visible state, explicit reasoning based from policies, strong intervention points, and rehearsal on realistic cases, grew more comfortable relying on the agent for both planning and execution. In two real-world deployments of the system, we found that organizers adopted DoubleAgents as transparency, policy controls, and edge‑case flagging built trust; however, required more upfront context setup and tighter email integration.
@@
Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic workflow. Specifically, we contribute:
  • A policy-driven, intervention‑friendly approach to govern proactive AI behavior in an agentic and integactive ReAct framework to build trust with users.
  • An AI-driven simulation framework of human responses to support repeated observations of reliable behavior over time to test and tune policies for AI to follow.
  • A user study (n=10) and a deployment with two users to evaluate how users perceive and gradually trust proactive AI for a human coordination task of scheduling a speaker seminar.
@@ 2.1. Agentic AI in Task Planning and Execution
“AI Agents” typically refers to AI that can perform tasks for people. This can be done through tool use (namely, calling an API) (Schick et al., 2023), or operating a computer (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025). Although AI Agents show impressive potential to act, trusting agents to act on behalf of a human is a critical missing piece (Afroogh et al., 2024). The delegation of control to AI raises questions not only about competence, but also about alignment, transparency, and user oversight.
Planning and execution remain central yet unresolved challenges in the development of AI agents. A growing body of work highlights both the difficulty and the importance of task planning, as seen in benchmarks and simulation environments such as Tau-Bench (Yao et al., 2024), NaturalPlan (Zheng et al., 2024), TravelPlanner (Xie et al., 2024), and AppWorld (Trivedi et al., 2024). These studies underscore that robust, goal-directed planning, particularly in open-ended or multi-step domains, is not only technically demanding but also essential for delivering meaningful user value.
To address this challenge, recent work has explored how LLMs can scaffold user planning by decomposing high-level goals into actionable steps (Zhang et al., 2024). Techniques such as ReAct (Yao et al., 2023b), which interleave reasoning and acting, have gained traction in multi-step environments for improving agent adaptability and responsiveness. These approaches are often embedded within multi-agent systems to support domains like travel, scheduling, and personal task management (Lal et al., 2023). However, autonomous systems such as AutoGPT (Significant Gravitas, [n. d.]) often fall short in transparency, adaptability, and robustness, limiting their usability in real-world settings (Zheng et al., 2024; Yao et al., 2024). Furthermore, prior cognitive science research shows that humans rarely engage in exhaustive upfront planning due to cognitive load (Krieger et al., 2009; Teevan et al., 2016), highlighting the need for AI systems that can support incremental, iterative, and contextualized planning over time. Additionally, recent studies suggest that human–AI co-planning, particularly when agents expose their intermediate reasoning and allow users to guide or intervene, can improve both task performance and user trust (He et al., 2025). When users can monitor and influence agent decisions at each step, they are better able to understand, calibrate, and ultimately trust the system, especially in domains requiring high-stakes delegation or multi-party coordination.
@@ 2.2. Human Simulations with LLMs
Recent work has explored the use of LLMs in multi-agent simulations  (Park et al., 2023), revealing their surprising capacity to model (Horton, 2023), predict, and even exhibit human-like behavior - including taking actions they were not explicitly instructed to perform in order to achieve high-level goals such as forming routines, organizing events, and collaborating spontaneously (Park et al., 2022, 2023). These behaviors often mirror human reasoning and coordination strategies. Other studies have shown that these behaviors closely align with empirical observations of humans across various domains, including economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023; Horton, 2023), U.S. Supreme Court judgments (Hamilton, 2023), and over 80% of published and unpublished psychology experiments (Hewitt et al., 2024). These results suggest that LLM agents can serve as realistic proxies for humans in testing coordination, negotiation, and reasoning behaviors at scale.
AI systems that silate human behavior have already shown a striking capability to engage in complex, social behaviors competitively well with humans, for real world purposes. . Beyond simulation, AI systems like CICERO (Team et al., 2022) have shown how language agents can engage directly with humans in dynamic, multi-party settings. In online Diplomacy games, CICERO achieved a top 10% ranking among human players - illustrating the potential for LLM-based systems to engage in trust-sensitive planning, social reasoning, and strategic communication. The MATRIX framework (Pang et al., 2024) improved alignment in constitutional AI evaluations by useing simulated AI dialogues to anticipate a range of perspectives and surface important value trade-offs. Many systems have successfully used LLM red teaming  (Purpura et al., 2025) to identify potential threats in systems. Although AI simulations might not perfectly replicate individual user behaviors (yet) - they can be used to explore edge cases, probe systems, and engage in sensitive dialogs with humans.
2.3. Constitutional AI
Constitutional AI is a framework for aligning AI systems with human values by training models to self-critique and revise their outputs based on a set of guiding principles, or a ”constitution”(Bai et al., 2022). Unlike reinforcement learning with human feedback (RLHF)(Ouyang et al., 2022), which requires human evaluation for each model response, Constitutional AI allows the model to evaluate its own behavior using pre-defined rules. These constitutions are typically expressed as human-readable policies such as “Don’t promote illegal activity” or “Be helpful and respectful,” and can be authored by anyone. Building on this idea, Collective Constitutional AI (Huang et al., 2024) introduced a participatory approach in which over 1,000 people contributed rules to a shared constitution. Models trained on this crowd-sourced constitution demonstrated qualitatively different behaviors compared to those trained on developer-authored rules—highlighting the diversity of human values and community norms, and the implications this has for model alignment. To create constitutions, interactive systems like ConstitutionMaker (Petridis et al., 2024) have shown that LLMs can help transform users’ natural language feedback into generalized alignment policies. By providing real-time AI assistance in policy creation, such systems help users articulate their values more effectively, reduce cognitive burden, and make the process of shaping AI behavior more accessible and inclusive. These advances open up opportunities for systems that incorporate user-driven policy creation and alignment into workflows that have real-world impacts.
3. Seminar Planning: A Human Coordination Problem
Human coordination—such as assigning work shifts or scheduling job interviews—is both cognitively demanding and socially delicate. Unlike tasks with stable inputs and predictable outcomes, coordination requires managing uncertainty, reconciling multiple perspectives, and sustaining relationships over time. Even small lapses—missed follow-ups, unclear tone, or misaligned expectations—can erode trust and derail collective progress. These challenges make coordination both essential and persistently difficult in many areas of knowledge work.
We study these challenges through one concrete and consequential setting: organizing an academic seminar series. A typical series spans an entire semester and involves inviting a new external speaker nearly every week, amounting to 8–12 speakers per term. Early on, organizers must identify potential speakers, gather their availability constraints, and balance their own, such as ensuring topical diversity across weeks and securing highly sought-after senior researchers who are harder to schedule.
@@ 4. DoubleAgents System
We present DoubleAgents, an interactive agentic system designed to investigate how users build and calibrate trust in proactive AI systems. Built on the ReAct framework (Yao et al., 2023b), DoubleAgents enables iterative planning and execution through a team of AI agents that assist users in complex coordination tasks. In this work, we focus on the use case of assigning seminar speakers to available time slots. The system takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots.”) and produces as output a finalized assignment of speakers to slots. DoubleAgents consists of six core components, grouped into the following three functional categories. First, it contains the Core ReAct Agents that carry out the reasoning, planning, and simulation logic that drive the coordination process:
@@
  • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions.
@@
  • Interactive Chat Interface - Facilitates communication between the user and the system. Users can confirm or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.

### v2

# DoubleAgents: Interactive Simulations for Alignment in Agentic AI

@@ Abstract.
Agentic workflows promise efficiency, but adoption hinges on whether people can align systems that act on their behalf with their goals, values, and situational expectations. We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks. A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their use of agentic behavior before live deployment. We evaluate DoubleAgents in a two-day lab study (n=10), three deployment studies, and a technical evaluation. Results show that participants initially hesitated to delegate but used simulation to probe system behavior and adjust policies, gradually increasing delegation as agent actions became better aligned with their intentions and context. Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that simulation helps users effectively manage real-world tasks with higher complexity and uncertainty. We contribute interactive simulation as a practical pathway for users to iteratively align and calibrate agentic systems.
@@
copyright: acmcopyrightjournalyear: 2025conference: 2026; 2026; arxivbooktitle: Proceedings of the 2026price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
Figure 1. DoubleAgents is an agentic AI coordination tool that uses interactive simulation to let users experience and align the full ReAct cycle before live deployment. Starting from an initial setup (left)—including a simulation goal, simulated test-case personas, and an initial set of policies, and email templates—users can undergo multiple simulated ReAct cycles (middle) to rehearse planning, review and approve actions, execute plans, detect new edge cases, and iteratively refine policies via stop hooks, and email templates over repeated test cases and time. This process results in (right) the final speaker-slot allocation, as well as a user-aligned set of policies, emails, and stop hooks before the system sends real emails.
@@ 1. Introduction
Human–AI alignment is one of the defining challenges of the 21st century, especially for agentic AI: systems that go beyond answering questions to taking actions on users’ behalf (Leike et al., 2018; Soares and Fallenstein, 2014; Hendrycks et al., 2021). Following the Bidirectional Human–AI Alignment framework defined by Shen et al. (2025), we view alignment in agentic settings as an ongoing co-adaptation process between users and systems. On the system side, agents must be continuously shaped through interaction history and curated context so their actions reflect an individual user’s goals, preferences, and social norms (Zhang et al., 2025; Gao et al., 2024). On the human side, users must continually develop appropriate mental models, usage patterns, and AI literacy so they can effectively direct, evaluate, and calibrate delegation (Vanderlyn et al., 2025; Andrews et al., 2023). As agentic AI becomes more common, we need practical ways to assess and improve whether these systems remain aligned as they act in the world.
Today, many popular agentic use cases, such as agentic coding or “vibe coding”, are relatively low-stakes and operate in closed sandbox environments (Horikawa et al., 2025). In these domains, alignment is often easier because systems can leverage large-scale, well-structured public data (e.g., GitHub), and success criteria are comparatively legible (e.g., whether code compiles, passes tests, or produces the desired output) (Chen et al., 2021; Jimenez et al., 2024). However, in more complex tasks that must adhere to human and domain knowledge as well as social cognition and norms, such as writing emails to real people to request favors, where etiquette is highly context-dependent, agents cannot reliably align on their own (Afroogh et al., 2024; Gabriel, 2020). These tasks often require nuanced, case-by-case judgment that is difficult to infer from generic training data alone and may demand substantial human oversight if deployed naively.
One representative scenario is human coordination, such as scheduling workers with the right skills and experience for shifts they are available for or inviting researchers to a speaker series. In this work, we focus on the speaker-series scheduling setting. Organizers must secure a speaker for each slot (e.g., 4-5pm on Fridays), coordinate availability across multiple candidates, and handle uneven responsiveness (e.g., some invitees reply late or not at all). Responses can also introduce socially sensitive and context-dependent requests: “Can you reimburse travel?”, “Can I do Zoom?”, “I can’t make it on Fridays, what about Thursdays 4-5pm?” “Sorry for the last-minute cancellation—I now have a conflict,” or “My schedule is uncertain; can you follow up next month?” Determining the right next step is rarely optimal in a mathematical sense; it depends on the organizer’s values and preferences, which we refer to as “policies” (e.g., “Offer Zoom only if in-person is not feasible and no one else can do that slot,” “Prioritize senior faculty over PhD students when conflicts happen,” or “Follow up after two business days unless urgent”). In practice, because these “actions” are carried out almost entirely through email communication, the organizer must decide which email to send next (and to whom), how to phrase each email with the right content and right tone, and when to send it.
While an LLM-based agent may generate plausible actions and write acceptable emails, it often lacks the organizer’s specific norms and cannot reliably anticipate how preferences should change across cases. Alignment therefore must happen on both sides: the system needs ways to acquire and apply three user-specific automation refinements: (1) coordination policies that guide planning and action choices, (2) email templates and writing style that match the user’s voice and social intent for each action, and (3) stop hooks that define when the agents need to pause and request for user confirmation; and the user needs support to form a clear mental model and specify these three preferences, especially early on, when they may not yet know what rules, tone, or escalation boundaries they want. The central challenge lies in how can we accelerate this bidirectional alignment so the agent can act more autonomously over time, without requiring constant babysitting.
Our core approach is interactive simulation to support iterative automation refinement, enabling alignment between users and agents. As modern AI can simulate diverse human behaviors (Park et al., 2023, 2022), we use simulation as a testbed to expose users to realistic coordination scenarios, including problematic or edge-case behaviors, and to help them define and refine their communication strategies. During simulations, the agent proposes plans and actions (e.g., inviting multiple candidates, deciding when to wait versus follow up, or holding a slot for a high-priority speaker), and the user evaluates whether these choices are socially and contextually appropriate. This mirrors ideas from software engineering such as unit and integration testing: a system earns confidence not by succeeding on a single happy path, but by handling a range of representative cases. In our context, however, completing a task is not enough: agentic AI systems must be user-specific and value-aligned, reflecting personal priorities, norms, and preferences.
We present DoubleAgents, a system for aligning agentic AI through interactive simulation. DoubleAgents integrates generative AI to simulate two complementary roles, as Figure 2 shows. First, it uses a ReAct loop (Yao et al., 2023) that pairs coordination agents for planning and execution. The agents can call tools to retrieve relevant context, query the current task state, and execute actions such as drafting and sending emails to potential speakers. We extend the loop with explicit interaction checkpoints so users can review, intervene, and approve decisions throughout planning and execution. Second, DoubleAgents includes an LLM-based simulator of respondents that generates diverse, realistic replies for different personas (e.g., a professor who responds slowly or a respondent who introduces funding or logistics constraints). This simulation enables repeated run-throughs of difficult cases, helping users understand how the agent behaves, stress-test policies against edge cases, and iteratively refine preferences and autonomy boundaries. Across interactive simulation runs, the primary outputs are not only a candidate schedule (speaker–slot allocation), but also a reusable set of user-specific alignment artifacts: refined policies, updated email templates, and calibrated stop hooks that can be applied when the system is deployed to send real emails.
We evaluate DoubleAgents through a technical evaluation, a two-day lab study (n=10) in which participants used DoubleAgents with simulated respondents, and three real-world deployments. In the lab study, participants initially hesitated to delegate. However, over time, through multiple simulation rounds and the two study sessions, participants increasingly delegated, as visible state, policy-grounded reasoning, and clear intervention points (including stop-hook flagging) helped them calibrate autonomy and build confidence in the agent’s planning and execution. By the end, participants also produced reusable policies, stop hooks, and email templates aligned with their voice and style. Three week-long deployment studies demonstrate DoubleAgents’ usefulness and relevance for real-world, complex human coordination tasks. DoubleAgents was connected to organizers’ real email workflows and sending emails to real invitees, as well as incorporated incoming real speakers’ replies into its ongoing coordination state. Organizers reported that interactive simulation beforehand helped them prepare for deployment by crafting strong initial policies and email templates for real-world use.
@@
Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic AI workflow. Specifically, we contribute:
  • A design perspective on alignment in agentic AI, realized through a policy-driven, intervention-friendly ReAct framework that supports transparency, human oversight, and alignment with human goals and values.
  • An interactive simulation mechanism that models realistic human responses, enabling users to safely explore agent behavior, understand system dynamics, rehearse interventions, and iteratively refine policies, email references, and stop hooks before real-world application.
  • Empirical results from a user study (n=10) and three real-world deployments, examining how interactive simulation and policy refinement support alignment in agentic AI for human coordination.
@@ 2. Related Works
2.1. Alignment in Human-AI Systems
As AI agents take on more complex tasks, ensuring they align with human values has become a central research challenge. A key strategy is reinforcement learning from human feedback (RLHF), which fine-tunes models to produce outputs consistent with user expectations (Ouyang et al., 2022). However, alignment doesn’t end at training. Studies show that users often push back on AI responses that conflict with their values, steering behavior through real-time feedback, a process termed user-driven value alignment (Fan et al., 2025). Such interactive alignment is critical in high-stakes contexts like medical decision support, where LLM-powered co-pilot systems have helped clinicians catch more errors than either humans or AI alone (Ong et al., 2025). These findings show the importance of transparency, oversight, and human agency in AI deployment.
Another approach, Constitutional AI, replaces constant human feedback with rule-based self-supervision. Models are trained to revise outputs using explicit principles like “Be respectful” (Bai et al., 2022). Recent work has expanded this into participatory frameworks such as Collective Constitutional AI, where large groups co-author alignment rules to reflect diverse values (Huang et al., 2024). Tools like ConstitutionMaker further help users turn natural language feedback into usable policies (Petridis et al., 2024). DoubleAgents contributes to this space by embedding alignment into interactive simulation. Users iteratively refine coordination policies, email tone, and stop hooks through rehearsed ReAct loops, treating alignment not as a one-time step, but as a continuous, participatory process.
@@ 2.2. Agentic AI in Task Planning and Execution
AI agents, systems that perform tasks on users’ behalf, operate through tool use (e.g., APIs) (Schick et al., 2023) or direct interface manipulation (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025). While these agents show growing capabilities, delegating control to them still raises concerns around trust, transparency, and alignment (Afroogh et al., 2024). Planning and execution remain core challenges. Benchmarks like Tau-Bench, NaturalPlan, TravelPlanner, and AppWorld (Yao et al., 2024; Zheng et al., 2024; Xie et al., 2024; Trivedi et al., 2024) highlight the difficulty of robust multi-step planning, especially in open-ended tasks. To address this, methods like task decomposition (Zhang et al., 2024) and ReAct, which interleaves reasoning and acting, have been shown to improve agent adaptability (Yao et al., 2023), especially when used in collaborative or multi-agent domains (Lal et al., 2023).
However, fully autonomous agents like AutoGPT often struggle with robustness and transparency in real-world use (Zheng et al., 2024; Yao et al., 2024). Cognitive research suggests humans rarely plan exhaustively upfront (Krieger et al., 2009; Teevan et al., 2016), pointing to a need for agents that support incremental and context-sensitive planning. Recent studies show that exposing agent reasoning and allowing user intervention improves both performance and trust (He et al., 2025). In sum, planning in agentic AI is not just a technical hurdle; it is a design problem. Transparent reasoning, iterative refinement, and human-in-the-loop oversight are essential for building trustworthy agents that can meaningfully support real-world tasks.
@@ 2.3. Human Simulations with LLMs
Recent work has shown that LLMs in multi-agent simulations can exhibit surprisingly human-like behaviors, forming routines, organizing events, and coordinating spontaneously without explicit instructions (Park et al., 2023; Horton, 2023; Park et al., 2022). These behaviors often reflect human reasoning and social strategies, aligning with patterns observed in economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023), judicial decisions (Hamilton, 2023), and over 80% of psychology experiments (Hewitt et al., 2024). This suggests LLM agents can serve as useful proxies for modeling human behavior at scale.
Beyond simulation, LLMs are demonstrating real-time social competence. CICERO, a dialogue-enabled Diplomacy agent, ranked in the top 10% of human players by integrating language with strategic planning (Team et al., 2022). Similarly, the MATRIX framework used simulated AI dialogues to surface moral tradeoffs in Constitutional AI evaluations (Pang et al., 2024), showing how LLMs can assist in value alignment. LLM simulations also support red teaming and adversarial testing, enabling safe, scalable evaluation of edge cases and vulnerabilities (Purpura et al., 2025). While imperfect, these agents provide a reproducible and low-risk method for probing system behavior in socially sensitive contexts, offering a powerful complement to human-centered alignment and evaluation.
@@ 3. DoubleAgents System
We present DoubleAgents, a system for aligning agentic AI through interactive simulation. Built on the ReAct framework (Yao et al., 2023), DoubleAgents supports iterative planning and execution via a team of AI agents that collaborate with users on complex coordination tasks. In this paper, we focus on a concrete use case: assigning seminar speakers to available time slots (see Appendix A for additional background). DoubleAgents takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots”), persona descriptions for simulated respondent agents, the user’s existing email templates, and a set of generic scheduling policies. Through iterative simulations, the system produces a finalized assignment of speakers to slots, along with refined policies, updated email templates, and executable stop hooks that users can later deploy in real-world agentic workflows. The system comprises six core components organized into three functional categories. First, DoubleAgents includes the Core ReAct Agents, which implement the reasoning, planning, and simulation logic that drives the coordination process:
@@
  • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions. These escalations are recorded as stop hooks, enabling iterative refinement and extension of the policy set.
@@
  • Interactive Chat Interface - Facilitates communication between the user and the system. Users can approve or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.